OpenStudy (anonymous):

Math Brain Twister: 100^2 - 99^2 + 98^2 - 97^2 + 96^2 - 95^2 + ... + 2^2 - 1^2 = ? Anyone? I'll give medal.. :))

hartnn (hartnn):

its like 100+99+98+97+....+2+1

hartnn (hartnn):

n(n+1)/2

random231 (random231):

sum of special series?

hartnn (hartnn):

n= 100

hartnn (hartnn):

5050

hartnn (hartnn):

no

hartnn (hartnn):

100^2 - 99^2 = (100+99)(100-99) = 100+99

random231 (random231):

hartnn is crrect @liliegirl

OpenStudy (anonymous):

Solution Since you can notice that there are fifty pairs of n^2 – (n-1) ^2, n^2 – (n-1)^2 = n + (n – 1) Thus 100^2 - 99^2 + 98^2 - 97^2 + 96^2 - 95^2 + ... + 2^2 - 1^2 can also be written as 100 + 99 + 98+ ... + 2 + 1 = (100 x 101)/2 = 5050

OpenStudy (anonymous):

Thanks! @hartnn

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

You guys are just using memory. Your explanations don't make sense to anyone who isn't familiar with the problem already.\[ n^2-(n-1)^2=n^2-(n^2-2n+1) = 2n-1 \]

OpenStudy (anonymous):

So if you're summing over \(2n-1\) when stepping by \(2\), how does that imply you are summing over \(n\) when stepping by \(1\)?